A variable line through the point P(2,1 ) meets the axes at A and B . Find the locus of the centroid of triangle OAB Where O is the origin 

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
Let the line =
y – 1 = m(x – 2)
 
y – 1 = mx – 2m
 
A (2m – 1/m , 0)
 
B (0, 1 – 2m)
 
Now 
Centroid = (2m – 1/ 3m , 1 – 2m/3)
 
Now you can solve it further.